A 1.9862-g sample of benzoic acid burns in an excess of oxygen, in a bomb calorimeter. The temperature of the calorimeter rises from 24.94 C to 28.77 C. The calorimeter contains 2000.0 g of H2O, and the heat capacity of the calorimeter (excluding the water, but including the bomb, thermometer, and all mechanical parts) has the value C(cal) = 5.35 kJ/C. Use the given information to find the molar heat of combustion of benzoic acid (M(benzoic acid) = 122.1 g/mol).
Answer #1

   2C6H5COOH(l) + 15O2(g) ----> 12CO2(g) + 6H2O(l)

no of mol of benzoicacid = w/mwt = 1.9862/122.1 = 0.163 mol

heat released(q) = mwater*swater*DT + Ccal*DT

                 = 2000*4.184*(28.77-24.94) + 5.35*10^3*(28.77-24.94)

                 = 52.54 kjoule

DH0com = -q/n = -52.54/0.163 = -322.33 kj/mol